A physics book slides off a horizontal tabletop with a speed of 1.10 m/s. It strikes
the floor in 0.350 s. Ignore air resistance. Find:
a. The height of the tabletop above the floor.
b. The horizontal and vertical components of the book’s velocity
1
Expert's answer
2020-08-24T11:56:30-0400

Given:

"v_{0x}=1.10\:\rm m/s"

"t=0.350\:\rm s"

(a) The height of the tabletop above the floor


"h=\frac{gt^2}{2}=\frac{9.81\times (0.350)^2}{2}=0.601\:\rm m"

(b) The horizontal component of the book’s velocity


"v_x=v_{0x}=1.10\:\rm m/s"

The vertival component of the book’s velocity

"v_y=gt=9.81\times 0.350=3.43\:\rm m/s"